A bag contains seven batteries, all of which are the same size and are equally likely to be selected. Each battery is different brand. If you select five batteries at random, use the counting principle to determine how many points will be in the sample space if the batteries are selected.
A) with replacement
b) without replacement

Answer by stanbon(75887)

A bag contains seven batteries, all of which are the same size and are equally likely to be selected. Each battery is different brand. If you select five batteries at random, use the counting principle to determine how many points will be in the sample space if the batteries are selected.
A) with replacement:::
7*7*7*7*7 = 7^5 = 16807
b) without replacement
Ans: 7*6*5*4*3 = 7P5 = 2520
=========================
Cheers,
Stan H.
=================